# JMLR: Joint Medical LLM and Retrieval Training for Enhancing Reasoning and Professional Question Answering Capability

### v2

@@ 2 Related Work
Medical LLMs have evolved from adapting existing models like BERT using biomedical datasets (Gu et al., 2021; Lee et al., 2020), to integrating document linking (Yasunaga et al., 2022b) and knowledge graphs (Yasunaga et al., 2022a). Decoder-only architectures such as GPT and Llama were then trained with domain-specific data (Radford et al., 2018; Touvron et al., 2023; Soldaini et al., 2023; Wu et al., 2023; Gema et al., 2023; Yunxiang et al., 2023; Zhang et al., 2023; Labrak et al., 2024) or some synthetic data (Tran et al., 2023; Han et al., 2023; Kweon et al., 2023; Wang et al., 2023b). Recently, scaling up data and model parameters led to models like GatorTronGPT (Peng et al., 2023) and Clinical-Camel (Toma et al., 2023), with studies on Flan-PaLM and PaLM-2’s medical reasoning (Singhal et al., 2023a, b). Furthermore, researchers have developed MEDITRON (Chen et al., 2023b), an open-source suite of medical LLMs with 7B and 70B parameters, offering enhanced performance compared to state-of-the-art baselines and outperforming closed-source models like GPT-3.5 and Med-PaLM, while closely rivaling GPT-4 and Med-PaLM-2 in effectiveness. Nonetheless, the persistent augmentation in the model’s size has not yielded commensurate enhancements in its performance metrics. This parameter expansion fails to adequately remedy the phenomenon of model forgetting (Wang et al., 2023a; Luo et al., 2023), and it appears that the model’s parameters are insufficient for retaining extensive knowledge. In contrast to existing approaches, we combine retriever and LLM in our training process to mitigate the forgetting issue. Our work provides two versions of medical LLMs: 7B and 13B. Evaluation results indicate that our models surpass previous traditional pre-trained language models in performance, and our 13B model even outperforms the MEDITRON 70B model.
@@
In the clinical domain, the occurrence of factual inaccuracies by language models (LMs), attributed to the gap in medical knowledge, presents considerable risk, with potentially severe implications such as misdiagnoses Petroni et al. (2019); Sung et al. (2021); Yao et al. (2022); Singhal et al. (2023a). Enhancing LMs by incorporating pertinent information sourced from diverse knowledge repositories has been demonstrated as an effective method for boosting performance across multiple Natural Language Processing tasks, notably in the realm of language modeling (Min et al., 2022). The retrieval component remains static in traditional RAG (Retrieval-Augmented Generation) models, while updates are applied solely to the LLM. Specifically, the input is utilized as a query in a two-step process. Initially, a retrieval system selects a collection of documents, which are essentially sequences of tokens, from a designated corpus. Subsequently, a language model assimilates these retrieved documents, employing them as supplementary information to formulate a conclusive prediction (Yu, 2022; Izacard et al., 2022). However, merely integrating the documents retrieved into the LM does not necessarily mean that the LM can use external knowledge. Consequently, several researchers have shifted the focus towards devising methods for guiding models in applying external knowledge. For instance, Atlas (Izacard et al., 2022) enhances an encoder-decoder model in tandem with the retriever by treating documents as latent variables. RETRO (Borgeaud et al., 2022b) modifies the decoder-only architecture to include retrieved texts and undertakes the pretraining of the LM from the ground up. However, LLMs are not inherently structured to seamlessly integrate retrieved content, often leading to discrepancies or inconsistencies when blending pre-trained model responses with external information. Fusing externally retrieved information with the model’s internal knowledge base can be intricate and flawed. This sometimes results in an awkward or inconsistent amalgamation of information, where the LLM may struggle to incorporate the external data into its responses coherently. Even though some researchers have considered training retriever on unsupervised data so that it can provide documents that are well-matched with LLM, such as REALM (Guu et al., 2020), REPLUG (Shi et al., 2023) and RA-DIT (Lin et al., 2023), these retriever systems are trained independently rather than in synchronization with LLM training. Our method integrates and synchronously trains the retriever and LLM, enabling the retriever to return documents more suited to LLM, thus enhancing LLM’s performance. Additionally, we have also conducted comparisons with step-by-step training methods to prove the effectiveness of our method.
3 Materials and Methods
There is a notable shortfall in biomedical knowledge datasets designed for training instruction-compliant assistants. To address this, we have collected the dataset of biomedical knowledge resources, primarily sourced from a wide array of clinical guidelines and textbook data, for pretraining. For the fine-tuning phase, we have gathered an extensive collection of medical QA datasets, aimed at augmenting the question-answering capabilities of LLMs.
3.1 Datasets
Knowledge Sources:
Data quality is important for retrieval of JMLR and for pretraining and finetuning LM. In routine practice, clinicians typically encounter at least one patient-care-related question for every two patients they see. They often turn to guidelines like UpToDate 222www.uptodate.com. UpToDate is an expert-authored, evidence-based clinical knowledge system that helps clinicians to make point-of-care decisions. Other high-quality resources include textbooks and electronic health record (EHR) notes. Our collection of UpToDate, textbooks and MIMIC-IV notes (the only publicly available EHR notes) has been used as documents to retrieve.
The fine-tuning process is based on the pre-trained Llama model, where we leveraged datasets from two significant sources: the USMLE and AMBOSS, which are essential for preparing for the USMLE examinations. The USMLE dataset, sourced directly from the official USMLE website, included a range of sample questions for Step1, Step2CK, and Step3, released between June 2022 and March 2023. These questions represent the complex medical knowledge and ethical scenarios that medical students and practitioners are expected to navigate. Additionally, we utilized the AMBOSS question bank, a comprehensive resource widely used by medical professionals and students. This dataset provided an extensive array of Step1, Step2CK, and Step3-type questions, further enriching our training material with practical and diverse medical scenarios. The MedMCQA dataset comprises 194k multiple-choice questions from Indian medical entrance exams, covering 2.4k healthcare topics and 21 medical subjects (Pal et al., 2022). The MMLU dataset Hendrycks et al. (2020), covering 57 subjects, including STEM and social sciences, features four-option multiple-choice questions for each subject. We focused on nine subjects most pertinent to medical and clinical knowledge — high school biology, college biology, college medicine, professional medicine, medical genetics, virology, clinical knowledge, nutrition, and anatomy.
3.2 Joint Medical LLM and Retrieval Training (JMLR)
3.2.1 LLM
The original Llama model supports a maximum of 4k tokens, which poses a challenge when multiple documents are retrieved. To address this, we utilized the Shifted Sparse Attention (S2-Attn) mechanism (Chen et al., 2023a), which can be used to process long sequences by mitigating the high memory cost and slow processing time associated with standard self-attention in LLMs. S2-Attn achieves this by dividing the input into several sequence spans within the self-attention modules and then introducing a shifted pattern to facilitate information exchange between these spans. This approach efficiently extends the context length without incurring additional computational costs and demonstrates performance close to the standard attention baseline. The method, designed for efficient training, retains the standard self-attention architecture during inference, making it compatible with existing infrastructure and optimizations.
3.2.2 Retriever
3.2.3 JMLR Architecture
However, Colbert matches queries and documents based on the retriever similarity, returning scores and documents. To enable the retriever system to return documents that are more helpful for providing accurate answers to Llama, we constructed the JMLR architecture. This structure employs a rank loss, , which links the Llama and retriever systems. The updating of retriever parameters is influenced by Llama’s content quality. If the returned document reduces Llama’s loss function more than document , it indicates that document is more helpful for answering than document . In calculating the rank loss, we incorporate the scores generated by the retriever system. However, due to the large variance in scores, we normalize them to obtain . The essence of is that the retriever system’s score rankings should be updated based on the ranking. in .
For JMLR model training, the dataset comprises QA pairs from AMBOSS and USMLE, with the input being a question and the output being the correct choice, including the reasoning part. The top 7 documents with the highest scores are aggregated and fed into the model during the training process. Prior to training, we extracted the top 30 documents based on scores from the retriever. During each iteration, 30 scores are generated by the retriever, and we perform weighted random sampling based on these scores, with higher scores having a greater probability of selection. This approach ensures that each document inputted into the LLM (Large Language Model) is dynamic. The LLM will likely encounter different useful guidelines and some that are not useful. This step enhances the LLM’s robustness, helping it discern which knowledge from which documents is useful and which documents are irrelevant.
3.3 Training Details
Both our training process and the conventional finetuning approach employ the AdamW optimizer, with , , and . We implement a cosine learning rate schedule, incorporating a warmup phase that accounts for 10% of the training duration and decays the learning rate to 10% of its peak value. In alignment with the practices outlined in Llama 2-chat touvron2023b, our training employs a learning rate of , a weight decay factor of 0.1, and manages a batch size of 2. The finetuning phase spans 5 epochs for all iterations. However, we apply a distinct learning rate for ColBERT, set at . The optimization strategy for training ColBERT mirrors that used for Llama 2. Throughout these experiments, we utilize four A100 80 GB GPUs. The total training duration for our method, when applied to the MedQA and AMBOSS datasets, amounts to 37 hours. Additionally, the pretraining phase on medical guidelines takes approximately 127 hours, while the finetuning phase is approximately 17 hours on 7B Llama. This comprehensive training approach ensures the effective adaptation and optimization of the model for specific medical applications.
4 Main Results
4.1 Effective of Retrieval
Setup: For benchmarks with available public training sets, we individually apply supervised finetuning using JMLR on each set and subsequently assess their performance on the corresponding test sets. PubMedQA and MedMCQA both offer detailed reasoning traces, such as extensive answers or explanations. For MedQA, AMBOSS, and MedMCQA, we employ an alternative training set equipped with human-crafted explanations for each query. In this scenario, JMLR is trained to attach the answer directly to the explanation. Regarding MMLU-Medical, a dataset without its own training set, we opt to evaluate it using the JMLR model originally trained on MedQA and AMBOSS. We calculate the accuracy of different medical QA datasets based on different models.
| Datasets | MMLU-Medical | MedMcQA | MedQA | AMBOSS |
| PMC-Llama-7B | 59.7 | 57.6 | 42.4 | 43.7 |
| Llama 2-7B | 56.3 | 54.4 | 44.0 | 46.5 |
| Meditron-7B | 55.6 | 59.2 | 47.9 | 50.1 |
| JMLR-7B | 57.2 | 61.3 | 51.7 | 68.7 |
Results: In Table 1, our JMLR-7B model, operating at the 7B scale, demonstrates its superiority by outperforming a variety of state-of-the-art (SOTA) public models across several benchmark datasets in the medical field. The datasets include MMLU-Medical, MedMcQA, MedQA, and AMBOSS. The enhanced performance of JMLR-7B can be partially attributed to its use of Llama-2 as the foundational model, showcasing significantly higher average performance than other pre-trained baselines. However, the distinct edge comes from our integrated training approach that combines retriever and language model learning (LLM). This methodology not only further boosts the performance of Llama-2 in medical benchmarks but also enables the retriever component to effectively assist the LLM in sourcing relevant and helpful documents for answering questions. This synergy significantly reduces instances of hallucination and enhances the overall effectiveness of the model. In the MMLU-Medical and MedMcQA datasets, JMLR-7B achieved scores of 57.2% and 61.3%, respectively, surpassing the scores of PMC-Llama-7B (59.7% and 57.6%), Llama 2-7B (56.3% and 54.4%), and Meditron-7B (55.6% and 59.2%). The trend continues in the MedQA and AMBOSS datasets, where JMLR-7B not only outperforms these models but does so with a notable margin, scoring 51.7% in MedQA and an impressive 68.7% in AMBOSS. This is particularly significant when compared to Meditron-7B, which scores 47.9% and 50.1% in these datasets, respectively. Overall, the results illustrate that the JMLR-7B model, on average, achieves about a 10% improvement in performance over its closest competitor, Meditron-7B. This highlights the effectiveness of our training method and the substantial advancements it brings to the field of medical benchmarking.
4.2 Effective of Joint LLM and Retrieval Training
Setup: In our experiment, we compared the 7B and 13B Llama models. We employed conventional RAG, fine-tuning the 7B and 13B models(RAG-7B and RAG-13B). During training, we didn’t update the parameters of the retriever; only the parameters of Llama were updated. We also separated the training steps: initially, we fixed the parameters of the retriever and only updated those of Llama, and then fixed Llama’s parameters while updating the retriever’s using rank loss, resulting in JLMR separate models, JLMR-separate-7B and JLMR-separate-13B. Finally, we trained JLMR using our integrated approach to obtain JMLR-7B and JMLR-13B. Additionally, we compared these models with larger-scale models such as ChatGPT and Meditron-70B.
Results: As shown in Figure 2, while the traditional RAG method offers some model accuracy enhancement, its impact is not particularly substantial. However, when we update the retriever using our rank loss(our-separate-7B and our-separate-13B), there is a noticeable improvement over the traditional RAG. Our-separate-7B and our-separate-13B achieved 51.8% and 65.4% in MedQA, and 54.5% and 71.3% in AMBOSS, respectively, outperforming the RAG variants. This improvement stems from the retriever’s enhanced ability, after the update, to identify documents that aid Llama in accurately answering questions. However, asynchronous training can be a bottleneck, especially in the efficient use of medical guidelines. While updating Llama’s parameters, the static nature of the retriever’s guidelines means that Llama might miss out on some crucial guidelines. This gap leads to the underutilization of potentially beneficial medical guidelines. Our joint training process,(JMLR-7B and JMLR-13B), effectively mitigates this issue, leading to a significant boost in performance. Because the parameters of the retriever are continuously updated during our training method, the best documents returned also change, allowing the model to encounter different documents that are helpful in answering questions, thereby enabling it to make better use of these medical guidelines. Compared to the larger models, JMLR-13B not only outperforms ChatGPT (which scores 55.4% in MedQA and 55.6% in AMBOSS) but also surpasses Meditron-70B. Specifically, JMLR-13B achieves an impressive 61.3% in MedQA and 81.2% in AMBOSS, indicating a substantial leap in model effectiveness and underscoring the success of our joint training approach.
4.3 Effective of Diverse Training Processes
Setup: To explore the impact of pretraining, finetuning, and information retrieval on model performance, we designed the following experiments: We tested the original Llama 2-7B on MedQA and AMBOSS datasets (Llama-7B). We also added a retriever during the inference stage to extract documents to assist Llama 2-7B in answering questions(Llama-7B+Retriever). Additionally, we finetuned Llama 2-7B on different training sets to obtain FT-Llama-7B. The same operation was repeated, adding a retriever during the inference stage to extract relevant documents to assist the model in answering questions(FT-Llama-7B+Retriever). Furthermore, to investigate the effectiveness of pretraining, we pretrained Llama 2 on medical guidelines(Llama-7B-Pretrain) and repeated the tests with and without adding a retriever (FT-Llama-7B-Pretrain, FT-Llama-7B-Pretrain+Retriever). Finally, we applied our training approach for further training on Llama-7B and Llama-7B-Pretrain (JMLR-7B, JMLR-7B-Pretrain).
Result: Figure 3 displays the varied performance of models at different stages. We observed that Llama-7B+Retriever, compared to Llama 2-7B, did not improve but decreased effectiveness. This suggests that Llama 2-7B does not utilize these guidelines effectively, and longer guidelines may even lead to misunderstandings and hallucinations. In contrast, FT-Llama-7B significantly improved over Llama-7B, indicating that the finetuning phase effectively enhanced Llama’s medical QA capabilities. However, this phase did not aid in the utilization of medical guidelines. FT-Llama-7B-Pretrain further improved upon FT-Llama-7B, and adding Retriever during the inference stage resulted in a slight increase in performance. This improvement is likely due to the pretraining phase expanding Llama 2-7B’s medical knowledge, thereby enabling it to utilize medical guidelines more effectively. Lastly, JMLR-7B and JMLR-7B-Pretrain achieved the best results. However, the minor difference between JMLR-7B and JMLR-7B-Pretrain suggests that our method has already sufficiently expanded Llama 2-7B’s medical knowledge. Considering training time, JMLR-7B required only 37 hours, while JMLR-7B-Pretrain needed 127 hours (for pretraining) plus 37 hours (for our method), indicating that our approach is more efficient in expanding the model’s medical knowledge.
4.4 Effective of the reasoning ability
Setup: To verify that our model can accurately answer questions and provide detailed reasoning, we tasked JMLR-13B with generating reasons for its responses to the AMBOSS and MedQA test sets. Similarly, we had ChatGPT generate corresponding reasons. We employed two methods for scoring these reasons: 1. We used GPT-4 to judge which model’s generated reasoning was better based on the ground truth. 2. We engaged three medical professionals to review and assess a sample of 20 questions drawn from the test sets. These doctors were asked to judge which model’s reasoning was closer to the actual, clinically accurate reasons.
Result: Table 2 demonstrates the varying win rates of our model under two different evaluation methods. In the assessment using GPT-4, JMLR-13B achieved a win rate of 0.63, while under the evaluation of medical professionals, our model attained a win rate of 0.60. These results suggest that simply increasing the model’s parameters is not the most effective way to expand medical knowledge. Instead, JMLR-13B’s ability to consult different medical guidelines for various questions and to find corresponding answers and explanations significantly reduces hallucinations and greatly enhances the model’s explanatory power. This achievement highlights the model’s proficiency in accessing and applying relevant medical information contextually, thereby providing more accurate and credible explanations. The close alignment of the win rates from both GPT-4 and medical professionals’ evaluations further underscores the model’s robustness and reliability in medical question-answering and reasoning tasks.
| Comparison | Win Rate |
| GPT4 | 0.63 |
| Expertise | 0.6 |
@@ 6 Limitations and Ethical Considerations
Despite significant advancements in the performance of medical knowledge acquisition and question-answering systems, privacy protection becomes a paramount concern when dealing with sensitive medical data. In particular, our model utilizes data from textbooks and Uptodate. Textbooks are public datasets, while the UpToDate dataset cannot be publicly disclosed or used to produce derivative products based on their terms of use. We have sent an email to the official authorities and received their permission to use it. We would like to express our gratitude to UpToDate and must ensure that no information about UpToDate is disclosed.
@@ A.1 Case Study
We presented four specific examples to further understand why JLMR outperforms other public models. The first three examples are from the AMBOSS dataset. Due to privacy concerns with AMBOSS, we cannot display the complete content of these questions. The last example is from a USMLE question. Our answers for the first two questions are correct, while ChatGPT responded incorrectly. Although both our model and ChatGPT answered the last two questions correctly, there were some differences in the reasoning generated. For the first example, we found that JLMR’s retrieval function could extract similar questions, leading the model to answer incorrectly regarding direction. The retrieved document served as a background, fully utilizing the previously fine-tuned dataset. In the second example, we noticed that ChatGPT made a common-sense error: vaccination is not recommended for pregnant women in such cases. Even though ChatGPT has learned far more medical documents than our model by using pretraining or finetuning, it does not mean that the model will apply all medical knowledge correctly, even for some basic medical issues. In the third example, both we and ChatGPT answered correctly, but the reasoning generated by ChatGPT differed significantly from the correct answer’s reasoning: ChatGPT simply stated some facts without providing a logical reason. However, our model gave a more detailed and logical explanation of why there are decreased circulating T cells, as JLMR could retrieve related documents, such as some documents about SCID, to better understand the underlying principles. The last example differs from the first three, as the highest-scoring document retrieved was from UpToDate’s guideline. The guideline provides a detailed introduction to PCOS and its symptoms, helping the model better explain. In contrast, ChatGPT simply correlated the symptoms straightforwardly without thoroughly analyzing other related symptoms, which can often lead to diagnostic errors, even though it answered this question correctly. Our model, however, performed a comprehensive analysis.
@@ A.2 Accuracy Variation with Different Numbers of Documents
ChatGPT: The correct answer is A) Cerebral edema. Explanation: This patient presents with symptoms consistent with alcohol withdrawal syndrome (AWS), which can manifest as seizures (alcohol withdrawal seizures) due to abrupt cessation or reduction of alcohol intake in a chronic heavy drinker. AWS seizures typically occur within 48 hours of reducing alcohol intake. This patient’s laboratory findings are indicative of hyponatremia (low serum sodium), which is common in severe alcohol withdrawal due to the syndrome of inappropriate antidiuretic hormone secretion (SIADH). SIADH leads to water retention and dilutional hyponatremia. The risk of cerebral edema is increased in patients with severe hyponatremia who are rapidly corrected to normal serum sodium levels. Rapid correction of hyponatremia can lead to osmotic shifts of water into the brain cells, causing cerebral edema, which may result in neurological deterioration, seizures, coma, and death. This risk is particularly high in chronic alcoholics due to adaptive changes in brain osmolytes. Given the patient’s current status, urgent treatment to normalize serum sodium levels (such as with hypertonic saline infusion) puts him at increased risk for cerebral edema. Therefore, careful monitoring and slow correction of hyponatremia are essential in this patient to avoid this potentially fatal complication. |
@@
Our: This patient’s presentation is consistent with alcohol withdrawal syndrome (AWS), which is characterized by autonomic hyperactivity, tremulousness, and seizures. AWS can occur in patients with a history of chronic alcohol use disorder who abruptly stop drinking or significantly reduce their alcohol intake. The syndrome is thought to be due to the withdrawal of the inhibitory effects of alcohol on the central nervous system. The most common complication of AWS is seizures, which can occur within 48 hours of the last drink. Seizures are often generalized tonic-clonic seizures and can be life-threatening. Other complications include delirium tremens (DTs), which is characterized by hallucinations, confusion, and autonomic instability. DTs can also lead to seizures and death. The treatment of AWS involves supportive care, including hydration, benzodiazepines, and possibly antipsychotics. In severe cases, hospitalization may be necessary for close monitoring and management of complications. Osmotic myelinolysis (also known as central pontine myelinolysis) is a condition that can occur in patients with severe hyponatremia (low sodium levels in the blood) who are treated with rapid correction of their sodium levels. This condition is characterized by demyelination of the central nervous system, leading to neurologic symptoms such as confusion, seizures, and coma. Given the patient’s presentation and history of chronic alcohol use disorder, the most appropriate treatment is supportive care for AWS, including hydration, benzodiazepines, and possibly antipsychotics. Rapid correction of his sodium levels is not indicated due to the risk of osmotic myelinolysis. |
@@
ChatGPT: Based on the clinical presentation described, the most likely diagnosis for this patient is Rubella (German measles) infection during pregnancy, which can lead to congenital rubella syndrome (CRS) in the fetus if contracted during the first trimester. The typical symptoms of rubella include a maculopapular rash starting on the face and spreading downwards, along with lymphadenopathy and joint swelling. Given the risk of congenital rubella syndrome to the fetus, the most appropriate next step in management is: C) Vaccination Vaccination against rubella is contraindicated during pregnancy due to the risk of causing CRS. However, after delivery, the patient should receive the rubella vaccine to prevent future infections. This will not only protect the patient from future rubella infections but also prevent the risk of transmitting the virus to others, including vulnerable populations such as pregnant women. Options A, D, and E are not appropriate in this scenario. Oral acyclovir therapy (Option A) is used for herpes simplex virus infections, not rubella. Reassurance and symptomatic therapy (Option D) may provide comfort but do not address the potential risk to the fetus. Intravenous immunoglobulin therapy (Option E) is not indicated for rubella infection during pregnancy. Termination of pregnancy (Option B) is not necessary unless there are severe fetal abnormalities detected on further evaluation. |
Our: The patient’s history of a rash that started on her face and progressed downwards, along with the presence of postauricular lymphadenopathy, is suggestive of rubella infection. Rubella is a highly contagious viral infection that is spread through respiratory droplets or direct contact with infected secretions. It is characterized by a prodrome of malaise, low-grade fever, and lymphadenopathy, followed by the appearance of a pink maculopapular rash that begins on the face and spreads downwards. The rash is usually itchy and lasts for 3-5 days. Rubella infection during pregnancy, especially during the first trimester, can have serious consequences for the developing fetus. It can cause congenital rubella syndrome (CRS), which is characterized by a wide range of birth defects, including cataracts, deafness, heart defects, and intellectual disability. The most appropriate next step in management for this patient is termination of pregnancy. This is because the risk of CRS is high if the fetus is exposed to rubella during the first trimester, and the risk of birth defects outweighs the potential benefits of continuing the pregnancy. Options A, B, C, and E are not appropriate in this case. Acyclovir is used to treat herpes simplex virus infections, not rubella. Vaccination is not recommended during pregnancy due to the risk of fetal infection. Reassurance and symptomatic therapy are not sufficient to prevent the risk of CRS. Intravenous immunoglobulin therapy is not indicated for rubella infection during pregnancy. |
@@
| 1 | Q3: A 1-year-old girl is brought to the physician… |

### v3

@@ 8 Related Work
Medical LLMs have advanced from adapting models like BERT with biomedical datasets (Gu et al., 2021; Lee et al., 2020) to incorporating knowledge graphs (Yasunaga et al., 2022). Subsequently, architectures such as GPT (Bubeck et al., 2023) and Llama (Touvron et al., 2023) were trained on domain-specific (Wu et al., 2023; Gema et al., 2023; Yunxiang et al., 2023; Zhang et al., 2023; Labrak et al., 2024) or synthetic data (Tran et al., 2023; Han et al., 2023; Kweon et al., 2023). Recent efforts have scaled up data and model parameters, resulting in GatorTronGPT (Peng et al., 2023) and Clinical-Camel (Toma et al., 2023), along with studies on Flan-PaLM and PaLM-2’s medical reasoning using chain-of-thought (Singhal et al., 2023a, b; Wei et al., 2022; Wang et al., 2022). Researchers have also introduced Meditron (Chen et al., 2023b), an open-source medical LLM suite with 7B and 70B parameters, demonstrating superior performance against leading benchmarks and comparable results to GPT-3.5 and Med-PaLM-2. Despite the model size increase, performance gains have plateaued, and such expansions have not effectively addressed model forgetting (Wang et al., 2023; Luo et al., 2023). In response, our study integrates a retriever with LLMs during training to tackle this issue, offering 7B and 13B versions. Our findings show that these models exceed traditional pre-trained language models in performance metrics, with our 13B model surpassing even the Meditron 70B.
@@
In the clinical domain, LLMs’ factual inaccuracies due to gaps in medical knowledge pose significant risks, including misdiagnoses Petroni et al. (2019); Sung et al. (2021); Yao et al. (2022); Singhal et al. (2023a). Incorporating diverse knowledge repositories into LLMs enhances their performance across various NLP tasks, particularly in language modeling (Min et al., 2022). In traditional RAG models, the retrieval component is static, with updates applied only to the LLM. The process involves using the input as a query to first retrieve a collection of documents from a designated corpus, which the LLM then uses to enhance its predictions (Yu, 2022; Izacard et al., 2022). However, integrating retrieved documents does not guarantee effective use of external knowledge. Innovations like Atlas (Izacard et al., 2022) and RETRO (Borgeaud et al., 2022b) have modified models to better utilize retrieved information, but challenges remain in coherently blending this external data with the LLM’s pre-trained responses. Although some researchers have explored unsupervised training of retrievers to improve document relevancy for LLMs, such as REALM (Guu et al., 2020), REPLUG (Shi et al., 2023), and RA-DIT (Lin et al., 2023), these systems are usually trained independently from LLMs. Our approach innovatively integrates and synchronously trains both the retriever and the LLM, ensuring the retriever supplies more appropriate documents, thereby boosting the LLM’s effectiveness. We have validated our method’s superiority through comparisons with step-by-step training techniques, confirming its enhanced performance in seamlessly integrating external knowledge.
@@ 10 Limitations and Ethical Considerations
Despite significant advancements in the performance of medical knowledge acquisition and question-answering systems, privacy protection becomes a paramount concern when dealing with sensitive medical data. In particular, our model utilizes data from textbooks and public guideline.
@@ A.1 Case Study
We presented four specific examples to further understand why JLMR outperforms other public models. The first three examples are from the Amboss dataset. Due to privacy concerns with Amboss, we cannot display the complete content of these questions. The last example is from a USMLE question. Our answers for the first two questions are correct, while ChatGPT responded incorrectly. Although both our model and ChatGPT answered the last two questions correctly, there were some differences in the reasoning generated. For the first example, we found that JLMR’s retrieval function could extract similar questions, leading the model to answer incorrectly regarding direction. The retrieved document served as a background, fully utilizing the previously fine-tuned dataset. In the second example, we noticed that ChatGPT made a common-sense error: vaccination is not recommended for pregnant women in such cases. Even though ChatGPT has learned far more medical documents than our model by using pretraining or finetuning, it does not mean that the model will apply all medical knowledge correctly, even for some basic medical issues. In the third example, both we and ChatGPT answered correctly, but the reasoning generated by ChatGPT differed significantly from the correct answer’s reasoning: ChatGPT simply stated some facts without providing a logical reason. However, our model gave a more detailed and logical explanation of why there are decreased circulating T cells, as JLMR could retrieve related documents, such as some documents about SCID, to better understand the underlying principles. The last example differs from the first three, as the highest-scoring document retrieved was from public guideline. The guideline provides a detailed introduction to PCOS and its symptoms, helping the model better explain. In contrast, ChatGPT simply correlated the symptoms straightforwardly without thoroughly analyzing other related symptoms, which can often lead to diagnostic errors, even though it answered this question correctly. Our model, however, performed a comprehensive analysis.
@@ A.2 Accuracy Variation with Different Numbers of Documents
| ChatGPT: The correct answer is A) Cerebral edema. Explanation: This patient presents with symptoms consistent with alcohol withdrawal syndrome (AWS), which can manifest as seizures (alcohol withdrawal seizures) due to abrupt cessation or reduction of alcohol intake in a chronic heavy drinker. AWS seizures typically occur within 48 hours of reducing alcohol intake. This patient’s laboratory findings are indicative of hyponatremia (low serum sodium), which is common in severe alcohol withdrawal due to the syndrome of inappropriate antidiuretic hormone secretion (SIADH). SIADH leads to water retention and dilutional hyponatremia. The risk of cerebral edema is increased in patients with severe hyponatremia who are rapidly corrected to normal serum sodium levels. Rapid correction of hyponatremia can lead to osmotic shifts of water into the brain cells, causing cerebral edema, which may result in neurological deterioration, seizures, coma, and death. This risk is particularly high in chronic alcoholics due to adaptive changes in brain osmolytes. Given the patient’s current status, urgent treatment to normalize serum sodium levels (such as with hypertonic saline infusion) puts him at increased risk for cerebral edema. Therefore, careful monitoring and slow correction of hyponatremia are essential in this patient to avoid this potentially fatal complication. | |
@@
| Our: This patient’s presentation is consistent with alcohol withdrawal syndrome (AWS), which is characterized by autonomic hyperactivity, tremulousness, and seizures. AWS can occur in patients with a history of chronic alcohol use disorder who abruptly stop drinking or significantly reduce their alcohol intake. The syndrome is thought to be due to the withdrawal of the inhibitory effects of alcohol on the central nervous system. The most common complication of AWS is seizures, which can occur within 48 hours of the last drink. Seizures are often generalized tonic-clonic seizures and can be life-threatening. Other complications include delirium tremens (DTs), which is characterized by hallucinations, confusion, and autonomic instability. DTs can also lead to seizures and death. The treatment of AWS involves supportive care, including hydration, benzodiazepines, and possibly antipsychotics. In severe cases, hospitalization may be necessary for close monitoring and management of complications. Osmotic myelinolysis (also known as central pontine myelinolysis) is a condition that can occur in patients with severe hyponatremia (low sodium levels in the blood) who are treated with rapid correction of their sodium levels. This condition is characterized by demyelination of the central nervous system, leading to neurologic symptoms such as confusion, seizures, and coma. Given the patient’s presentation and history of chronic alcohol use disorder, the most appropriate treatment is supportive care for AWS, including hydration, benzodiazepines, and possibly antipsychotics. Rapid correction of his sodium levels is not indicated due to the risk of osmotic myelinolysis. | |
@@
| ChatGPT: Based on the clinical presentation described, the most likely diagnosis for this patient is Rubella (German measles) infection during pregnancy, which can lead to congenital rubella syndrome (CRS) in the fetus if contracted during the first trimester. The typical symptoms of rubella include a maculopapular rash starting on the face and spreading downwards, along with lymphadenopathy and joint swelling. Given the risk of congenital rubella syndrome to the fetus, the most appropriate next step in management is: C) Vaccination Vaccination against rubella is contraindicated during pregnancy due to the risk of causing CRS. However, after delivery, the patient should receive the rubella vaccine to prevent future infections. This will not only protect the patient from future rubella infections but also prevent the risk of transmitting the virus to others, including vulnerable populations such as pregnant women. Options A, D, and E are not appropriate in this scenario. Oral acyclovir therapy (Option A) is used for herpes simplex virus infections, not rubella. Reassurance and symptomatic therapy (Option D) may provide comfort but do not address the potential risk to the fetus. Intravenous immunoglobulin therapy (Option E) is not indicated for rubella infection during pregnancy. Termination of pregnancy (Option B) is not necessary unless there are severe fetal abnormalities detected on further evaluation. | |
| Our: The patient’s history of a rash that started on her face and progressed downwards, along with the presence of postauricular lymphadenopathy, is suggestive of rubella infection. Rubella is a highly contagious viral infection that is spread through respiratory droplets or direct contact with infected secretions. It is characterized by a prodrome of malaise, low-grade fever, and lymphadenopathy, followed by the appearance of a pink maculopapular rash that begins on the face and spreads downwards. The rash is usually itchy and lasts for 3-5 days. Rubella infection during pregnancy, especially during the first trimester, can have serious consequences for the developing fetus. It can cause congenital rubella syndrome (CRS), which is characterized by a wide range of birth defects, including cataracts, deafness, heart defects, and intellectual disability. The most appropriate next step in management for this patient is termination of pregnancy. This is because the risk of CRS is high if the fetus is exposed to rubella during the first trimester, and the risk of birth defects outweighs the potential benefits of continuing the pregnancy. Options A, B, C, and E are not appropriate in this case. Acyclovir is used to treat herpes simplex virus infections, not rubella. Vaccination is not recommended during pregnancy due to the risk of fetal infection. Reassurance and symptomatic therapy are not sufficient to prevent the risk of CRS. Intravenous immunoglobulin therapy is not indicated for rubella infection during pregnancy. | |
@@
| 1 | Q4: A 1-year-old girl is brought to the physician… |